write a system of two linear equations in two variables for:a plane flew with a tail wind from madrid to lima, a trip of 1054 km in 93 min, the return trip took 9 more mins.
Let p be the air speed of the plane in km/h and w for the wind speed in km/h.
solve the system and find the air speed of the plane

Answer by mananth(16946)

speed of plane = p kmph
speed of wind = w kmph
..
speed with wind = p+w
speed against wind = p-w
..
distance = 1054 km
..
time with wind = 93 minutes = 1.55 hours
time against wind = 102 minutes= 1.7 hours
..
t=d/r
1054/p+w= 1.55
1054 =1.55p+1.55w...............1
...
1054/(p-w)=1.7
1054 = 1.7p+1.7w................2
...
solve these equations (1) & (2) to get the value of p & w
p= 650 km/h
w = 30 km/h
...
m.ananth@hotmail.ca